OpenStudy (anonymous):

Mrs. Horn made chocolate chip cookies for a party. She checks the recipe and sees that she needs 2 1/4 cups of flour for each batch. A batch makes approximately 3 dozen cookies. (a) How much flour would be needed for 7 batches of cookies? Show your work (b) How many batches of cookies would she need if she wanted to make at least 200 cookies to take to the party? Show your work (c) What is the unit rate and what does the number mean?

OpenStudy (anonymous):

\[2.25 \times 7 \] \[= 15.75 \] answer for a \[12 \times 3\] \[= 36\] \[5.5 \times 2.25 \] \[= 13.75 cups \] answer for b I don't know the other one